What is the specific heat (in J/(gram*oC)) of a substance if it takes 5400. Joules to heat 35g from 15 oC to 50. oC? Remember to apply significant figures. Note that some numbers have decimal points.
1
Expert's answer
2018-03-12T09:45:18-0400
Answer:
Specific heat is defined as the amount of heat per unit mass needed to increase the temperature by
one degree Celsius (or by 1 Kelvin).
Q = mcΔT 
Rearrange and solve for C:
C = Q/m ΔT  =5400J/35g x (50 – 15) oC =5400/ 1225 = 4.4 J/(gram*oC